OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) =48 when x=8, find the value of f(x) when x=2

OpenStudy (anonymous):

f(x)=kx 48=8k find k

OpenStudy (anonymous):

f(x) = kx k = f(x)/x f(8) = 48 k = 48/8 = 6 f(2) = 6. 2 = 12 so 12 is d answer